# ReaDy-Go: Real-to-Sim Dynamic 3D Gaussian Splatting Simulation for Environment-Specific Visual Navigation with Moving Obstacles

### v1

@@ Abstract
Visual navigation models often struggle in real-world dynamic environments due to limited robustness to the sim-to-real gap and the difficulty of training policies tailored to target deployment environments (e.g., households, restaurants, and factories). Although real-to-sim navigation simulation using 3D Gaussian Splatting (GS) can mitigate this gap, prior works have assumed only static scenes or unrealistic dynamic obstacles, despite the importance of safe navigation in dynamic environments. To address these issues, we propose ReaDy-Go, a novel real-to-sim simulation pipeline that synthesizes photorealistic dynamic scenarios for target environments. ReaDy-Go generates photorealistic navigation datasets for dynamic environments by combining a reconstructed static GS scene with dynamic human GS obstacles, and trains policies robust to both the sim-to-real gap and moving obstacles. The pipeline consists of three components: (1) a dynamic GS simulator that integrates scene GS with a human animation module, enabling the insertion of animatable human GS avatars and the synthesis of plausible human motions from 2D trajectories, (2) navigation dataset generation for dynamic environments that leverages the simulator, a robot expert planner designed for dynamic GS representations, and a human planner, and (3) policy learning using the generated datasets. ReaDy-Go outperforms baselines across target environments in both simulation and real-world experiments, demonstrating improved navigation performance even after sim-to-real transfer and in the presence of moving obstacles. Moreover, zero-shot sim-to-real deployment in an unseen environment indicates its generalization potential. Project page: https://syeon-yoo.github.io/ready-go-site/.
@@ I Introduction
Visual navigation policies using only an RGB camera provide practical advantages such as reduced robot hardware complexity, lower sensor cost, and lighter payloads. Furthermore, it is beneficial to integrate them with other RGB-based perception and decision-making modules in robotic systems. Despite these advantages, their performance in real-world dynamic environments often lacks robustness and efficiency.
There are several reasons for the limited visual navigation performance in the real world. First, the sim-to-real gap leads to a substantial drop in navigation performance. RGB-only navigation models commonly rely on deep learning to learn nonlinear visuomotor mappings from high-dimensional observations under monocular depth ambiguity. They are usually trained in simulation and deployed in the real world, as collecting large amounts of real-world training data is not feasible. Although the simulation data is essential for visual navigation policies, the distribution gap between simulation and real-world domains leads the policies struggle during robot deployment. Second, another cause of low performance is the difficulty of collecting datasets from target deployment environments and training environment-specific navigation policies for targeted scenes. Given that most real-world robots are operated in site-specific environments, such as households, restaurants, or factories, navigation policies tailored to the deployment environments are necessary for effective and robust performance. However, generating digital twins to collect datasets is expensive and labor-intensive [24]. On the other hand, although there have been attempts to train general navigation models [28, 29, 30], they still require a pre-built topological map. Also, navigation policies trained with large datasets show lower success rates than environment-specific policies in target environments, since general policies cannot fully leverage the specialized information of the deployed scenes [5, 19].
@@
To overcome these limitations, previous works have proposed a real-to-sim simulation using 3D Gaussian Splatting (GS) [11] for visual navigation policy training. These approaches mitigate the sim-to-real gap by using datasets generated from reconstructed scenes [5, 37, 22, 18, 20, 32, 21]. GS provides highly photorealistic 3D scene reconstruction, novel view synthesis, fast rendering, and explicit scene representations, which are beneficial for simulation and planning. Prior works leverage these advantages by reconstructing environments from RGB videos and generating photorealistic navigation datasets. Although a GS-based real-to-sim simulation is a promising approach for robust visual navigation, existing works assume only static environments or unrealistic dynamic obstacles based on simulation assets. However, addressing safe navigation in dynamic environments is essential for robots to be deployed in real-world scenarios. The generation of photorealistic dynamic-environment datasets for navigation remains underexplored.
@@
To improve visual navigation performance in real-world dynamic environments, we propose ReaDy-Go, a photorealistic Real-to-Sim Dynamic 3D Gaussian Splatting Simulation pipeline for environment-specific visual navigation with moving obstacles (Fig. 1). With the proposed pipeline, we can generate photorealistic navigation datasets for dynamic environments, and facilitate robust sim-to-real transfer of RGB-only visual navigation policies trained on the generated datasets. The key ideas are as follows: (1) building a dynamic GS simulator utilizing a static scene GS, dynamic human GS obstacles, and a human motion generation module; (2) generating photorealistic dynamic-environment datasets via a pipeline composed of the dynamic simulator, a robot expert planner designed for dynamic GS representations, and a human planner; and (3) training an RGB-only visual navigation policy using imitation learning with the generated datasets. We consider goal navigation as the downstream task, as it is a fundamental building block that can be extended to other navigation tasks. To the best of our knowledge, ReaDy-Go is the first pipeline that constructs a GS-based dynamic environment simulation for navigation dataset generation with animated human GS obstacles and trains navigation policies using the generated datasets.
In summary, our main contributions are threefold.
@@
  • We develop a photorealistic real-to-sim dynamic 3D Gaussian Splatting simulator with dynamic human GS obstacles. The simulator represents both the scene and the human as GS representations and generates human motion, i.e., body root locations and joint configurations, from a given 2D trajectory.
@@
  • We propose a new pipeline, ReaDy-Go, that generates photorealistic dynamic scenario datasets and trains environment-specific visual navigation policies, without cumbersome processes such as mesh extraction and physics engine integration. The pipeline incorporates the dynamic GS simulator, a robot expert planner designed for dynamic GS representations, and a human planner.
@@
  • The proposed method achieves robust RGB-only visual navigation performance in dynamic environments in both simulation and sim-to-real transfer experiments. Furthermore, it shows generalization via zero-shot sim-to-real deployment in an unseen environment.
@@ II-A RGB-Only Visual Navigation Policies
RGB-only navigation often uses learning-based methods to learn nonlinear visuomotor policies that map high-dimensional observations to actions while mitigating monocular depth ambiguity. Since learning-based navigation policies usually require a large amount of data, previous works have typically used simulation datasets to train visual navigation policies by imitation learning or reinforcement learning [31, 16, 25, 19, 34, 33]. To leverage the fact that simulation can easily capture multi-modal data, cross-modal learning that uses heterogeneous data during training while employing only RGB for inference has also been proposed to distill multi-modal information into RGB inputs [35, 2]. However, most RGB-only navigation policies struggle when transferring policies trained in simulation to the real world, although many of them attempt to mitigate the visual sim-to-real gap using methods such as domain randomization [16, 25].
Some studies have proposed methods to address the sim-to-real gap for visual navigation policies. One such line of work is General Navigation Models (GNMs), general-purpose navigation policies trained on the large-scale real-world navigation datasets [30, 29, 28]. Although GNMs show the potential to handle diverse robot embodiments and environments in a zero-shot manner using a single model, policies trained with large-scale data may not effectively harness the environment-specific information, leading to lower performance than environment-specific policies [5, 19]. In addition, their application is limited due to the requirement of a pre-built topological map, i.e., a sequence of goal images along a desired path. Another approach to mitigate the sim-to-real gap is to use a real-to-sim simulation that reconstructs real-world environments in a simulator to generate photorealistic navigation datasets. This method is useful for learning environment-specific navigation models that are robust and efficient in target environments. Real-to-sim simulation reconstructs the scene using asset retrieval [14, 6] or 3D Gaussian Splatting (GS). Our method leverages GS-based real-to-sim simulation, as it provides highly photorealistic and fast rendering, and explicit representations that are favorable for planning and control in simulation.
@@ II-B GS-Based Real-to-Sim Simulation for Navigation Policies
GS-based real-to-sim simulation is a promising research direction in the robotics field to reduce the sim-to-real gap for robot policies in navigation [5, 37, 22, 18, 20, 32, 21] and manipulation [23, 36, 17]. In navigation tasks, rule-based planning pipelines using a GS map [4, 13] have been proposed to leverage the benefits of GS representations, such as consistent geometry and high-fidelity rendering at high frame rates. However, they cannot respond to dynamic environments and require a GS map during inference, which can be difficult to employ on resource-limited robots. Concurrently, learning-based navigation policies trained in real-to-sim GS simulations have been presented across various hardware systems such as drones [22, 18, 20, 21], wheel-based robots [5, 32], and legged robots [37]. Real-to-sim simulations enable sim-to-real transfer of the navigation policies to target deployment environments and also show generalization potential of policies in novel scenes. However, they assume static environments, and set relatively easy tasks or strong goal cues, such as following predefined trajectories or finding red/blue circles.
@@
For real-world applications, considering dynamic environments is essential, but most previous GS-based real-to-sim simulation research does not. Although Vid2Sim [32] used simulation assets as dynamic obstacles, it requires cumbersome mesh extraction and physics engine integration processes to generate observations from heterogeneous representations, GS and mesh, and produces less photorealistic results, which may not fully leverage the benefits of real-to-sim. In contrast, our method, ReaDy-Go, develops a pipeline for navigation policies that is composed of a photorealistic real-to-sim dynamic GS simulation using GS for both the scene and dynamic human obstacles, together with a robot expert planner and a human planner. We show that ReaDy-Go improves sim-to-real transfer performance of environment-specific navigation policies in dynamic environments and demonstrates policy generalization.
@@ III-A1 Static GS scene reconstruction
The background scene in which the robot will be deployed is reconstructed using GS. GS is an explicit representation that enables 3D geometry reconstruction, high-fidelity novel view synthesis, and fast training and rendering by fitting positions, rotations, scales, opacities, and colors of 3D Gaussian primitives to the training set images [11]. Specifically, we employed PGSR [3] for 3D scene reconstruction, which achieves high-quality surface reconstruction and rendering by compressing 3D Gaussians into flat planes and using geometric regularization loss terms in addition to a photometric loss. Improved geometric accuracy compared to the vanilla GS reduces less Gaussian noise in the scene and improves multi-view consistency. This is important when extracting scene voxels and 2D occupancy grids for planners of the robot and dynamic obstacles.
@@ III-A2 Human Animation Module
Subsequently, to generate natural human motion along desired 2D trajectories of dynamic obstacles, PriorMDM [27] is adopted to predict the body root trajectory and joint angles in the SMPL parameters, which are used to animate human GS models. It enables fine-grained trajectory-level control over human motion using a pretrained motion diffusion model (MDM) as a generative prior. Given a 2D desired trajectory, we convert it into body root linear and rotation velocities, normalize them to match the HumanML3D [8] representation as the model input, and feed them into PriorMDM. The predicted 3D body joint positions are then fitted to SMPL parameters through SMPLify [1] and transformed into the world coordinate frame in the target environment, which makes plausible human animation along a desired trajectory possible.
@@ IV-B1 Dataset
We selected three target environments, Outside, Lobby, and Library, as shown in Figs. 4 (simulation) and 5 (real-world). Each environment was reconstructed as a GS scene using a monocular video recorded for about six minutes, which corresponds to 1,000–1,500 images. During ReaDy-Go simulation, we generated 400 training episodes for each environment, which corresponds to approximately 80k–120k data samples. The validation scenarios consist of 50 episodes for each environment, and we selected the checkpoint with the best validation performance, which is used for testing in simulation and the real world.
@@ IV-C Visualization of ReaDy-Go Dynamic GS Simulation
Qualitative results of the proposed dynamic GS simulation, ReaDy-Go, show that it generates photorealistic real-to-sim datasets for dynamic environments, as depicted in Fig. 4. The qualitative results across various target environments indicate that the simulation can animate human GS models in the target environments with plausible body motion along the given 2D trajectories without a physics engine. ReaDy-Go integrates photorealistic scene GS and human GS representations and is able to create thousands of navigation scenarios for dynamic environments using our human animation module, expert planner, and human planner. Furthermore, the novel view synthesis results demonstrate that the simulation can render navigation scenarios from arbitrary viewpoints with geometrically consistent photorealistic dynamic scenes.
@@ IV-D Visual Navigation Performance in Simulation
The effect of photorealistic dynamic GS simulation datasets for visual navigation policies is examined through simulation tests comparing ReaDy-Go against baselines, as depicted in Table I. We observe two key takeaways from the experiments. First, training environment-specific navigation policies using real-to-sim simulation is crucial for safe and efficient navigation performance. ReaDy-Go and Vid2Sim, which are trained in real-to-sim target environments, achieve higher success rates and lower average reaching times than general navigation models (GNM, NoMaD, and ViNT) across all tasks and environments, despite the general models requiring up to $15\times$ more parameters. Real-to-sim simulation with GS is a cost-effective and scalable approach to achieve fewer collisions and faster task completion with only a video.
Second, photorealistic dynamic obstacles in our simulation are a key factor in maintaining visual navigation performance in dynamic environments. The success rate and average reaching time of ReaDy-Go and Vid2Sim in Static are competitive across all environments. However, in Dynamic, ReaDy-Go shows higher success rates and lower average reaching times than the baselines, whereas Vid2Sim experiences a sharp drop. Since the two methods differ only in the training data, i.e., photorealistic human GS dynamic obstacles for ReaDy-Go versus human assets in a physics engine for Vid2Sim, these results suggest that the proposed photorealistic dynamic GS simulation helps improve navigation performance in dynamic scenarios, which is a practical advantage for robot deployment.

### v2

@@ Abstract
Visual navigation models often struggle in real-world dynamic environments due to limited robustness to the sim-to-real gap and the difficulty of training policies tailored to target deployment environments (e.g., households, restaurants, and factories). Although real-to-sim navigation simulation using 3D Gaussian Splatting (GS) can mitigate these challenges, prior GS-based works have considered only static scenes or non-photorealistic human obstacles built from simulator assets, despite the importance of safe navigation in dynamic environments. To address these issues, we propose ReaDy-Go, a novel real-to-sim simulation pipeline that synthesizes photorealistic dynamic scenarios in target environments by augmenting a reconstructed static GS scene with dynamic human GS obstacles, and trains navigation policies using the generated datasets. The pipeline provides three key contributions: (1) a dynamic GS simulator that integrates static scene GS with a human animation module, enabling the insertion of animatable human GS avatars and the synthesis of plausible human motions from 2D trajectories, (2) a navigation dataset generation framework that leverages the simulator along with a robot expert planner designed for dynamic GS representations and a human planner, and (3) robust navigation policies to both the sim-to-real gap and moving obstacles. The proposed simulator generates thousands of photorealistic navigation scenarios with animatable human GS avatars from arbitrary viewpoints. ReaDy-Go outperforms baselines across target environments in both simulation and real-world experiments, demonstrating improved navigation performance even after sim-to-real transfer and in the presence of moving obstacles. Moreover, zero-shot sim-to-real deployment in an unseen environment indicates its generalization potential. Project page: https://syeon-yoo.github.io/ready-go-site/.
@@ I Introduction
Visual navigation policies that rely solely on an RGB camera provide practical advantages for robotic systems, including reduced hardware complexity, lower sensor costs, and lighter payloads. Furthermore, RGB-only navigation modules can be seamlessly integrated with other vision-based perception and decision-making components. Despite these benefits, their performance in real-world dynamic environments often lacks robustness and efficiency, limiting reliable deployment in practical settings.
The fundamental challenge lies in achieving robust navigation in environment-specific, dynamic real-world settings. RGB-only navigation models typically learn nonlinear visuomotor policies from high-dimensional monocular observations, where depth ambiguity complicates scene understanding. Therefore, most learning-based approaches are trained extensively in simulation, as collecting large-scale real-world navigation data is impractical. However, the resulting sim-to-real distribution gap significantly degrades performance during deployment. Moreover, robots are commonly operated in environment-specific settings such as households, restaurants, or factories, where scene layouts are unique. General navigation models [28, 29, 30] often fail to fully exploit environment-specific characteristics, which leads to reduced success rates [5, 19], while constructing digital twins for targeted dataset generation remains expensive [24].
@@
To mitigate these limitations, recent works [5, 37, 22, 18, 20, 32, 21] have proposed real-to-sim simulation pipelines based on 3D Gaussian Splatting (GS) [11]. By reconstructing environments from RGB videos, GS enables high-fidelity rendering at fast frame rates, novel view synthesis, and simulation with an explicit 3D scene representation. GS-based pipelines leverage these advantages for navigation dataset generation and policy training to reduce the sim-to-real gap in target environments. However, existing GS-based navigation frameworks are limited to static scenes or introduce dynamic elements using non-photorealistic simulation assets, such as pre-defined human meshes. Such limitations make it difficult to learn safe navigation in the presence of dynamic obstacles and to render photorealistic human appearances within reconstructed real-world environments. As a result, the generation of photorealistic navigation datasets for dynamic environments remains underexplored.
Addressing this issue is essential for robot deployment in real-world environments with moving elements. In particular, there is currently no GS-based real-to-sim pipeline that jointly (1) models animatable human GS obstacles within reconstructed static GS scenes, (2) generates navigation datasets tailored to environment-specific training in dynamic settings, and (3) trains visual navigation policies that are robust under sim-to-real transfer and to moving obstacles.
@@
Motivated by these limitations, we propose ReaDy-Go, a photorealistic Real-to-Sim Dynamic 3D Gaussian Splatting Simulation pipeline for environment-specific RGB-only visual navigation with moving obstacles (Fig. 1). The proposed framework consists of three key components: (1) a dynamic GS simulator that integrates a static scene GS, an animatable human GS obstacle, and a human motion generation module, enabling the placement of a human in the scene and the synthesis of plausible motions conditioned on 2D trajectories; (2) a photorealistic dataset generation pipeline for dynamic environments that is composed of the dynamic simulator, a robot expert planner designed for dynamic GS representations, and a human planner; and (3) training an RGB-only visual navigation policy using imitation learning with the generated datasets. We focus on goal navigation as the downstream task, as it is a fundamental building block extensible to broader navigation problems.
To the best of our knowledge, ReaDy-Go is the first GS-based real-to-sim framework for dynamic scenes that enables photorealistic dataset generation with animated human GS obstacles and environment-specific navigation policy training. Our contributions are threefold.
@@
  • Dynamic GS Simulator: We develop a photorealistic real-to-sim dynamic 3D Gaussian Splatting simulator with human GS obstacles. The simulator represents both the scene and the human as GS representations and generates human motion, i.e., body root locations and joint configurations, from a given 2D trajectory.
@@
  • Photorealistic Dynamic Dataset Generation Pipeline: We propose a new pipeline, ReaDy-Go, that generates photorealistic dynamic scenario datasets and trains environment-specific visual navigation policies, without requiring mesh extraction and physics engine integration. The pipeline incorporates the dynamic GS simulator, a robot expert planner designed for dynamic GS representations, and a human planner.
@@
  • Robust Sim-to-Real Performance in Dynamic Environments: The proposed method achieves robust RGB-only visual navigation performance in dynamic environments in both simulation and sim-to-real transfer experiments. Furthermore, it shows generalization via zero-shot sim-to-real deployment in an unseen environment.
@@ II-A RGB-Only Visual Navigation Policies
RGB-only navigation often uses learning-based methods to learn nonlinear visuomotor policies that map high-dimensional observations to actions while mitigating monocular depth ambiguity. Since learning-based navigation policies usually require large amounts of data, previous works have used simulation datasets to train policies by imitation learning or reinforcement learning [31, 16, 25, 19, 34, 33]. To leverage the fact that simulation can easily capture multi-modal data, cross-modal learning that uses heterogeneous data during training while employing only RGB for inference has also been proposed to distill multi-modal information into RGB-only policies [35, 2]. Despite these advances, policies trained in simulation often suffer from sim-to-real performance degradation. Techniques such as domain randomization [16, 25] may not fully eliminate distribution gaps.
An alternative direction to address the sim-to-real gap for visual navigation policies is to train General Navigation Models (GNMs) on large-scale real-world navigation datasets [30, 29, 28]. GNMs show the potential to handle diverse robot embodiments and environments in a zero-shot manner using a single model. However, they typically require a pre-built topological map (e.g., sequences of goal images) during inference and may underperform compared to environment-specific policies in target deployment settings [5, 19], as they cannot fully exploit scene-specific structural information.
To address these limitations, real-to-sim simulation approaches reconstruct target environments and generate photorealistic datasets tailored for environment-specific policy training. Scene reconstructions can be achieved using asset retrieval [14, 6] or 3D Gaussian Splatting, which provides explicit scene representations for rendering and planning.
@@ II-B GS-Based Real-to-Sim Simulation for Navigation Policies
Real-to-sim simulation based on 3D Gaussian Splatting (GS) has recently gained attention in robotics for reducing the sim-to-real gap in both navigation [5, 37, 22, 18, 20, 32, 21] and manipulation tasks [23, 36, 17]. GS offers high-fidelity rendering at fast frame rates, novel view synthesis, and geometrically interpretable primitives that are useful for simulation.
For navigation, the first type of work is rule-based planning methods using GS maps to exploit their geometric consistency and rendering quality [4, 13]. While effective in static environments, these methods require access to the GS map at inference time and are not well suited for resource-constrained robotic platforms. In the second type of work, learning-based navigation policies are trained in GS-based real-to-sim environments and transferred to actual robot hardware such as drones [22, 18, 20, 21], wheeled robots [5, 32], and legged robots [37]. These approaches demonstrate improved sim-to-real transfer by leveraging photorealistic reconstructions of deployment environments.
Despite this advantage, existing GS-based real-to-sim navigation frameworks predominantly assume static scenes, whereas dynamic environments are essential for real-world applications. Moreover, some works simplify tasks by using predefined trajectories or strong goal cues (e.g., colored targets), reducing the task difficulty. Although Vid2Sim [32] introduces dynamic obstacles via simulation assets, it may limit photorealism compared to fully GS-based pipelines and requires additional mesh extraction and physics engine integration to generate observations from heterogeneous representations (GS and meshes).
@@
In contrast, our method, ReaDy-Go, develops a pipeline for navigation policies that is composed of a photorealistic real-to-sim dynamic GS simulation using GS for both the scene and dynamic human obstacles, together with a robot expert planner and a human planner. We show that ReaDy-Go improves sim-to-real transfer performance of environment-specific navigation policies in dynamic environments and demonstrates policy generalization.
@@ III-A1 Static GS scene reconstruction
The background scene in which the robot will be deployed is reconstructed using GS. GS is a representation that enables 3D geometry reconstruction, high-fidelity novel view synthesis, and fast training and rendering by fitting positions, rotations, scales, opacities, and colors of 3D Gaussian primitives to the training set images [11]. Specifically, we employed PGSR [3] for 3D scene reconstruction, which achieves high-quality surface reconstruction and rendering by compressing 3D Gaussians into flat planes and using geometric regularization loss terms in addition to a photometric loss. Improved geometric accuracy compared to the vanilla GS reduces Gaussian noise in the scene and improves multi-view consistency. This is important when extracting scene voxels and 2D occupancy grids for planners of the robot and dynamic obstacles.
@@ III-A2 Human Animation Module
Subsequently, to generate natural human motion along desired 2D trajectories of dynamic obstacles, PriorMDM [27] is adopted to predict the body root trajectory and joint angles in the SMPL parameters, which are used to animate human GS models. It enables fine-grained trajectory-level control over human motion using a motion diffusion model (MDM) as a generative prior. Given a 2D trajectory, we convert it into body root linear and rotation velocities, normalize them to match the HumanML3D [8] representation as the model input, and feed them into PriorMDM. The predicted 3D body joint positions are then fitted to SMPL parameters through SMPLify [1] and transformed into the world coordinate frame in the target environment, which makes plausible human animation along a desired trajectory possible.
@@ IV-B1 Dataset
We selected three target environments, Outside, Lobby, and Library, as shown in Figs. 4 (simulation) and 5 (real-world). Each environment was reconstructed as a GS scene using a monocular video recorded for about six minutes, yielding 1,000–1,500 images. We used six human GS avatars (Section III-A2) into the navigation datasets. During ReaDy-Go simulation, we generated 400 training episodes for each environment, which corresponds to approximately 80k–120k data samples. The validation scenarios consist of 50 episodes for each environment, and we selected the checkpoint with the best validation performance, which is used for testing in simulation and the real world.
@@ IV-C Visualization of ReaDy-Go Dynamic GS Simulation
Qualitative results demonstrate that ReaDy-Go produces photorealistic real-to-sim dynamic environments, as shown in Fig. 4. First, the proposed human animation module generates plausible body motions for human GS avatars within static GS scenes along given 2D trajectories, without relying on a physics engine. Second, the framework supports scalable generation of thousands of photorealistic navigation scenarios via expert and human planners. Finally, novel view synthesis results confirm that geometric consistency can be preserved from arbitrary viewpoints in dynamic settings.
@@ IV-D Visual Navigation Performance in Simulation
The impact of photorealistic dynamic GS simulation datasets on visual navigation policies is examined through simulation tests comparing ReaDy-Go against baselines, as summarized in Table I. We observe two key takeaways from the experiments. First, training environment-specific navigation policies using real-to-sim simulation is crucial for safe and efficient navigation. ReaDy-Go and Vid2Sim, both trained in real-to-sim target environments, achieve higher success rates and lower average reaching times than general navigation models (GNM, NoMaD, and ViNT) across all tasks and environments. This performance is particularly notable given that the general models requiring up to $15\times$ more parameters. These indicate that real-to-sim simulation with GS is a cost-effective and scalable approach to achieve fewer collisions and faster task completion with only a video.
Second, photorealistic dynamic obstacles in our simulation are a key factor in maintaining visual navigation performance in dynamic environments. While ReaDy-Go and Vid2Sim perform comparably in static environments, their performance differs significantly in dynamic settings. ReaDy-Go maintains high success rates and low average reaching times relatively well, whereas Vid2Sim exhibits a substantial performance drop. Since the two methods differ only in the training data, i.e., photorealistic human GS dynamic obstacles for ReaDy-Go versus human assets in a physics engine for Vid2Sim, these results suggest that the proposed photorealistic dynamic GS simulation helps improve navigation performance in dynamic scenarios, which is a practical advantage for robot deployment.